Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 621     Entry time: 18 Jun 2025, 09:47
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 18.06.2025 
General Announcement:
-  LIMP-25 information: Martin Doering informs: 
1. Monday 7AM the tunnels are open but smoke extraction system tests is planned from 8AM to 10AM so that tunnels will be noisy and doors difficult to open; 
2. Men's toilettes at the XHEXP1 will be closed for maintenance and the woman's will be for dual use; 
3. HXS construction activities start next week but respecting the grace period for the first two week (no infra-structure works); 
4. New screens are installed at the tunnels entrances with info of major activities taking place at each tunnels sections; 
5. Last LIMP planning meeting on Friday 20.06 with group's representatives and news will be shared during the XO Integration Meeting. 

CTRL:
- Due to issues with Broker the plan is to upgrade MID and HED topics in coming weeks. Timeline will be shared.

DA:
- All fine at DA.

EEE:
- Performing timing tests documented in XFEL e-log; XO informs the system was rolled back to the trigger with 50Hz mains yesterday by 17:00 due to tuning activities.

VAC:
- Nothing to report.

XPD:
- FD studies ongoing this week.

DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- A. Trapp mentions XRO completed run of temperature tests at M2 XTD10 as planned.
- No info on watchdog tests.

LAS:
- Bringing up Laser back after 50Hz tests

UNSYS:
- Nothing to report.

SPB:
- Nothing to report.

FXE:
- D3 tests are ongoing and seems to be well working.

MID:
- No issues to report.

HED:
- Not present.

SCS:
- Nothing to report.

SQS:
- D3 tests ongoing this week.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00